OpenStudy (anonymous):

Determine a Cartesian equation for the curve given in parametric form by x(t) = 2e^2t y(t) = 3e^−t.

OpenStudy (anonymous):

1. x^2y=6 2.x/y^2=12 3.xy^2=6 4.x^2/y=6 5.xy^2=18 6.x/y^2=18 *answer choices*

OpenStudy (dumbcow):

x/2 = e^(2t) sqrt(x/2) = e^t y = 3/e^t = 3/sqrt(x/2) --> y^2 = 9/x/2 = 18/x